OpenStudy (anonymous):

Evaluate the following expression for a = 4 and b = -5. a2 - 10 2 + b -8 6 8 -2

OpenStudy (anonymous):

so what is the question exactly? are there two separate equations?

OpenStudy (anonymous):

16-10= 6 (for the first problem and 2-5= -3 (second problem)